ROTATIONAL SPECTROSCOPY OF THE CCCCl RADICAL

\underline{TAKASHI YOSHIKAWA}, YOSHIHIRO SUMIYOSHI, and YASUKI ENDO, Graduate School of Arts and Sciences, The University of Tokyo.

Pure rotaional spectra of the CCCCl radical in a supersonic jet have been observed for the first time by FTMW spectroscopy. The radical was produced in a pulsed electric discharge of C_{2}H_{2} and CCl_{4} diluted to 0.6 and 0.3 % with Ar, respectively. Transitions with spin splittings and hyperfine splittings were observed for the two isotopomers, CCC^{35}Cl and CCC^{37}Cl, in the region from 11.3 GHz for N = 2-1 to 33.8 GHz for N = 6-5. The molecular constants including the hyperfine constants due to the Cl nucleus have been precisely determined.